OpenStudy (anonymous):

Can anyone please explain The Binomial Theorem and Pascal’s Triangle to me? I will medal?!

OpenStudy (anonymous):

Practice problems would be great too because I have no idea how to do it

ganeshie8 (ganeshie8):

First of all, recall that a "binomial" is a polynomial with "two" terms. so the theorem must have something to do with "two terms".

OpenStudy (anonymous):

Ok:)

ganeshie8 (ganeshie8):

Indeed it does, the binomial theorem gives a nice expression for repeated multiplication of a binomial with itself: \[(x+y)(x+y)(x+y)\cdots\]

OpenStudy (anonymous):

I see, I see

ganeshie8 (ganeshie8):

Without beating around the bush, here is the theorem in all its glory : \[\large (x+y)^n = \sum\limits_{k=0}^n \color{red}{\binom{n}{k}}x^ky^{n-k}\]

OpenStudy (anonymous):

I'm just gonna go ahead and do pascals triangle. In very simple words, its a number triangle. To build the triangle, start with "1" at the top, then continue placing numbers below it in a triangular pattern. Each number is the two numbers above it added together (except for the edges, which are all "1").

OpenStudy (anonymous):

:O Stop! The formula was NOT what I was expecting...

OpenStudy (anonymous):

But how does it relate to the binomial theorem

ganeshie8 (ganeshie8):

\(\large \color{red}{\binom{n}{k}}\) is called "binomial coefficient", pascal's triangle helps you in finding this coefficeint.

OpenStudy (anonymous):

Let’s explore the first couple of powers of the binomial (a + b). (a + b)0 = 1 *Remember, anything to the zero power equals 1. (a + b)1 = a + b (a + b)2 = (a + b)(a + b) = a^2 + 2ab + b^2 (a + b)3 = (a + b)(a2 + 2ab + b2) = a^3 + 3a^2b + 3ab^2 + b^3

OpenStudy (anonymous):

uh-huh I'm processing this slowly, but I feel like I'm getting it

OpenStudy (anonymous):

Now take a look at the pascals triangle. (just search it up online)

OpenStudy (anonymous):

compare the coefficients, and powers with the pascals triangle

OpenStudy (anonymous):

1 a + b a^2 + 2ab + b^2 a^3 + 3a^2b + 3ab^2 + b^3

OpenStudy (anonymous):

and 1 11 121 1331

OpenStudy (anonymous):

Yes, I understand that.

OpenStudy (anonymous):

Can any of you guys give me like a practice question to help maybe?

OpenStudy (anonymous):

The row represents the coefficients

OpenStudy (anonymous):

Hold on ffggff i just need to say one more thing

OpenStudy (anonymous):

kk:)

OpenStudy (anonymous):

Take a look at the exponents for (a+b)^3 The powers of a decrease from 3 to 0, and the powers of b increase from 0 to 3.

OpenStudy (anonymous):

so that is the relationship between the two Here is a practice problem Expand (x + 2)^5 using the Binomial Theorem and Pascal’s triangle.

OpenStudy (anonymous):

Side note, is that the fifth row of pascals triangle would represent (a+b)^4, and the third would be (a+b) ^2. The reason is that the first row, represents (a+b)^0

OpenStudy (anonymous):

Ok, I'm lost lol. So I drew the little triangle thingy and got 1, 5, 10, 10, 5, 1. But now what?

OpenStudy (anonymous):

Like I can do the triangle but I don't know what to do afterwards?

OpenStudy (anonymous):

The sixth row of Pascal’s triangle would be 1, 5, 10, 10, 5, 1 and the exponents would start and end with 5. And that is what you got

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

So then \[1x^5 2^0 + 5x^4 2^1 + 10x^3 2^2 + 10x^2 2^3 + 5x^12^4 + 1x^02^5 =\] \[x^5 + 5 • 2x^4 + 10 • 4x^3 + 10 • 8x^2 + 5 • 16x + 32 =\] \[x^5 + 10x^4 + 40x^3 + 80x^2 + 80x + 32\]

OpenStudy (anonymous):

:O Whoa! Okay! How did you add the 6th row of the triangle to the equation?

OpenStudy (anonymous):

I mean, how did you get those exponents?

OpenStudy (anonymous):

Remember this? Take a look at the exponents for (a+b)^3 The powers of a decrease from 3 to 0, and the powers of b increase from 0 to 3. SO basically, as it goes, the powers of x decrease from the n to 0 (where n is the exponent, or (x+y)^n) And the powers of the other one (in this case it is 2) increases from 0 to n.

OpenStudy (anonymous):

Ok :) Thank-you, I get it now!!!!

OpenStudy (anonymous):

NP!!! :)

OpenStudy (anonymous):

So basically they are coming from (a+b)^5 in a way?

OpenStudy (anonymous):

@zzr0ck3r I feel so bad! Lol you have been typing forever :)

OpenStudy (anonymous):

well, yes, So to finalize it let the equation be (a+b)^n So a's exponents would go from n to 0 and b's would increase from 0 to n.

OpenStudy (anonymous):

how do you what to do to break it up into 5? Like how the first one is 1x^5 2^0 and the second one is 5x^42^1?

OpenStudy (zzr0ck3r):

A neat trick you can use instead of the triangle. Example: \((a+b)^4\) We know that if will look something like this \(\text{_}a^4+\text{_}a^3b+\text{_}a^2b^2+\text{_}ab^3+\text{_}b^4\) Now the question is, what do we put in the _ places Well we know the first coefficient will be \(1\). For the second _ we multiply the power of \(a\) in the first term, by the coefficient in the first term and then divide that all by \(1\) (we use one because the previous two steps were in the first term). So \((4*1)/1=4\) and that is our second coefficient. so we have \(a^4+4a^3b+\text{_}a^2b^2+\text{_}ab^3+\text{_}b^4\) For the third coefficient, we multiply the power of \(a\) in the second term by the coefficient in the second term and divide by \(2\) (we are doing stuff in the second term now) and we get \(3*4/2 = 6\) and that is our third coefficient. Similarly our fourth coefficient is \(2*6/3=4\) and our last coefficient is \(1*4/4=1\) So we have \(a^4+4a^3b+6a^2b^2+4ab^3+b^4\). Also I am sure someone pointed this out above, but once we find the first half of the coefficients, we are done (notice the pattern). Ok this all seems very hard and long, but that was just because I was doing lots of steps. If wanted to do \((a+b)^7\) First we quickly write out \(a^7+a^6b+a^5b^2+a^4b^3+a^3b^4+a^2b^5+ab^6+b^7\) and note the leading coefficient is \(1\). then we do \(7*1/1=7\\ 6*7/2=21\\ 5*21/3=5*7=35\) and we are done because we have gone half way so we have \((a+b)^7=a^7+7a^6b+21a^5b^2+35a^4b^3+35a^3b^4+21a^2b^5+7ab^6+b^7\)

OpenStudy (anonymous):

do you see what I mean kinda?

OpenStudy (zzr0ck3r):

I did not realize it was going to be so long once I started...

OpenStudy (anonymous):

@zzr0ck3r Thank-you :), I might just use your way from now on. It's simple yet very effective!

OpenStudy (anonymous):

well yes @zzr0ck3r but ffggfgf wants to know the relationship between binomial theorem and pascals triangle.

OpenStudy (zzr0ck3r):

this actually shows it:)

OpenStudy (anonymous):

It's a good method.

OpenStudy (zzr0ck3r):

I can generally do these things and write them down faster than someone can put it in a calculator and write it down.

OpenStudy (anonymous):

i use a different method than these, it is quite fast also.

OpenStudy (zzr0ck3r):

But anyway, it was meant to be a little extra add-on to what you all were saying...

OpenStudy (anonymous):

i would write it, but i don't want to spend another hour lol

OpenStudy (anonymous):

Thanks to everyone who helped!! I really understand how to do it now!! :)

ganeshie8 (ganeshie8):

I think that works because of combinations, which term we pick in each binomial :